A ten-wheeler truck of mass 4500kg and traveling at 10 m/s hits the back of a parked car. The car and its occupants have a total mass of 960kg. What is the speed of the car and the truck immediately after impact assuming they are locked together? If the impact lasts for 0.4s, what force is experienced by a 75kg passenger


1
Expert's answer
2021-10-14T11:32:11-0400

Apply momentum conservation to determine the speed after the impact:


"Mv=(m+M)u,\\\space\\
u=v\frac{M}{m+M}=8.24\text{ m/s}."

The force is


"F=m_pa=m_p\frac ut=1545\text{ N}."